The distance between jfk airport and la is 2500 miles. Let s be airspeed of jet. Wind speed 100 miles.
A. Equation S took 2 he and 5 min From NY to la against wind. I need an equation for this.
B. Solve equation
C. Write and find out how long it takes to fly against wind between my and la. Wind =150 per hour and airspeed is 525 per hour.

Answer by mananth(16949)

d=2500
s= speed of jet
wind speed = 100 mph
time = 2 hours 5 mins = 2 5/60 = 2(1/12)= 25/12

25/12 = 2500/(x-100)
c. 2500/375